OpenStudy (anonymous):

A candy store sells malted milk balls for $2.50 per pound, and chocolate-covered raisins for $3.75 per pound. How many pounds of each should be used to make a 10 pound mixture that sells for $3.00 per pound?

OpenStudy (anonymous):

Set up two equations. One finds out "how much stuff", and one finds out "the cost". x is milk balls, y is raisins. x+y=10 lbs 2.50x+3.75y=3.00

OpenStudy (anonymous):

we have 2.5x+3.75y=30 x + y = 10 solve the system we have x=6 y = 4 6 pounds of malted milk balls and 4 pounds of covered raisins. Yum!

OpenStudy (anonymous):

Now use the substitution method. Set the first equation equal to y (or x). x=10-y plug in 10-y wherever u see an x in the second equation 2.50(10-y)+3.75y=3.00

OpenStudy (anonymous):

Distribute the 2.50 and solve for y. 25-10y+3.75y=3.00 25-6.25y=3.00 -6.25y=-22 y=3.52

OpenStudy (anonymous):

hey gandalf, your setup of the second equation has a mistake

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

it's 2.5x+3.75y=30

OpenStudy (anonymous):

because it's a ten-pound mixture and costs 3.00 per pound, and thus the total cost is going to be 30?

OpenStudy (anonymous):

OHHHH- missed that! Thank u! So, do the same steps except make 3 30.

OpenStudy (anonymous):

i got confused

OpenStudy (anonymous):

u want me 2 walk u through it lil?

OpenStudy (anonymous):

please, i got lost with the 3.30

OpenStudy (anonymous):

where did that come from?

OpenStudy (anonymous):

The candy-person wanted a 10 pound mixture costing 3 dollars per pound. Because there are ten pounds total, that means that the whole thing will cost 30 dollars. I had made a mistake and thought that the whole thing would only cost 3 dollars. I had meant in my former comment, if that's what confused you, that where I wrote 3.00, u should put down 30.00- sorry if I wrote it weird

OpenStudy (anonymous):

oh, ok

OpenStudy (gw2011):

Let M = Malted Milk Let C = Chocolate 2.50M + 3.75C = 30 M + C = 10 Solve for both equations by multiplying the second by 2.50 2.50M + 3.75C = 30 2.50M + 2.50C = 25 Subtract this equation from the first one 1.25C = 5 C = 4 Substitute this value into one of the equations to find M: M + C = 10 M + 4 = 10 M = 6 You need 4 pounds of Chocolate and 6 pounds of Malted Milk